Shift Summary for eve shift, Sept 22, 1999

Patricia McBride (mcbride@fnpsph)
Thu, 23 Sep 1999 00:08:38 CDT

Shift Summary for eve shift, Sept 22, 1999

On Shift: P. McBride, J. Graham

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included):0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: no beam, no spill signal; Tevatron studies

========================================================================
Runs ended this shift:
none

========================================================================
Accesses during shift:
Valeri and Patty made an access to adjust voltages on VV'.

BobT and RickF made an access into NM2

========================================================================
Overview of shift:
No beam; Jason, Ed, Margaret, Vince worked on the upgrade of the memories;
ended up returning to the initial configuration. They will work on it again
tomorrow during the shutdown. Jason tried to test DA with cosmics (no spill
signal).

Valeri adjusted voltages on VV'.

BobT and RickF modified the illumination of the accidental telescope
to reduce the rate

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

Test DA with beam or cosmics when spill signal returns.

========================================================================
Items to Follow up on Next Shift

-- 
Patricia McBride          mcbride@fnal.gov          PHONE: (630)840-8071
Fermi National Accelerator Laboratory               FAX:   (630)840-2783